OpenStudy (anonymous):

What is the simplified answer of √140

OpenStudy (anonymous):

11.83

OpenStudy (anonymous):

Simplified? You can split up one root into two roots. For example. \[\sqrt{140} = \sqrt{4} \times \sqrt{35}\] \[2\sqrt{35}\]